# Algorithmic contiguity from low-degree conjecture and applications in correlated random graphs

### v2

@@ 4 Detection in correlated SBMs
To prove Theorem 4.1, it suffices to show the following lemma, as incorporated in Subsection B.4 of the appendix.
@@
Acknowledgment. The author thanks Jian Ding and Jingqiu Ding for helpful discussions. The author thanks Alexander S. Wein for helpful comments on the revised low-degree conjecture. The author thanks Hang Du for pointing him to the reference [SW25]. This work is partially supported by National Key R&\&&D program of China (No. 2023YFA1010103) and NSFC Key Program (Project No. 12231002).
@@ Appendix B Supplementary proofs
${}M=M_{n}=\max\big{\{}n,\epsilon_{n}^{-\frac{1}{2}}\big{\}}\to\infty\,.$ (B.3)
@@
For “natural” high-dimensional hypothesis testing problems between $\mathbb{P}$ and $\mathbb{Q}$, the following statements are equivalent:
@@
    $\mathsf{Adv}_{\leq D}\big{(}\tfrac{\mathrm{d}\mathbb{P}^{\prime}}{\mathrm{d}\mathbb{Q}^{\prime}}\big{)}=O(1)$ (respectively, $\mathsf{Adv}_{\leq D}\big{(}\tfrac{\mathrm{d}\mathbb{P}^{\prime}}{\mathrm{d}\mathbb{Q}^{\prime}}\big{)}=1+o(1)$) as $n\to\infty$ for some $\mathbb{P}^{\prime},\mathbb{Q}^{\prime}$ such that $\operatorname{TV}(\mathbb{P},\mathbb{P}^{\prime}),\operatorname{TV}(\mathbb{Q},\mathbb{Q}^{\prime})=o(1)$.
@@
    There is no algorithm with running time $n^{\widetilde{O}(D)}$ that achieves strong detection (respectively, weak detection) between $\mathbb{P}$ and $\mathbb{Q}$.
@@
Assuming Conjecture C.1, we have for all $1\leq i\leq n$ and all statistics $\{g_{i,j}=g_{i,j}(A,B):1\leq j\leq n\}$ such that each $g_{i,j}(A,B)$ can be computed in time $n^{\widetilde{O}(D)}$ and
@@
Without losing of generality, we may assume that $i=1$ in the following proof. Suppose on the contrary that there are statistics $\{g_{j}=g_{j}(A,B):1\leq j\leq n\}$ that can be computed in time $n^{\widetilde{O}(D)}$ such that $\sum_{1\leq j\leq n}g_{j}(A,B)=o(n)$ for all graphs $A,B$ and (3.6) holds. Again we may assume that $0\leq g_{i}\leq 1$. Recall (3.7), (3.8) and (3.9). Using Lemma 3.5 and Theorem 2.4, we see that (assuming Conjecture C.1) we have (3.10) still holds. To this end, define $g=g_{1}+\ldots+g_{n}$. Using $0\leq g\leq n$, we see that (3.10) yields that $\mathbb{Q}(g\geq\Omega(n))\geq\Omega(n)$, which violates our assumption (C.1). This implies that (3.6) is impossible and thus completes our proof. ∎
@@
Suppose on the contrary there are statistics $\{f_{j}:1\leq j\leq n\}$ that can be computed in time $n^{\widetilde{O}(D)}$ satisfying Items (1) and (2) in Definition 3.1 with $\mathbb{E}_{\mathbb{P}_{*}}[f_{\pi_{*}(i)}]\geq 1-c$ for some fixed constant $0.010<c<$ 0.01. We first claim that we may assume that

### v3

@@ 4 Detection in correlated SBMs
To prove Theorem 4.1, it suffices to show the following lemma, as incorporated in Section B.4 of the appendix.
@@
Acknowledgment. The author thanks Jian Ding and Jingqiu Ding for helpful discussions, Alexander S. Wein for helpful comments on the revised low-degree conjecture, and Hang Du for pointing him to the reference [SW25]. This work is partially supported by National Key R&\&&D program of China (No. 2023YFA1010103) and NSFC Key Program (Project No. 12231002).
@@ Appendix B Supplementary proofs
${}M=M_{n}=\max\big{\{}n,\lceil\epsilon_{n}^{-\frac{1}{2}}\rceil\big{\}}\to\infty\,.$ (B.3)
@@
For “natural” high-dimensional hypothesis testing problems between $\mathbb{P}$ and $\mathbb{Q}$, the following statements hold.
@@
    If $\mathsf{Adv}_{\leq D}\big{(}\tfrac{\mathrm{d}\mathbb{P}^{\prime}}{\mathrm{d}\mathbb{Q}^{\prime}}\big{)}=O(1)$ as $n\to\infty$ for some $\mathbb{P}^{\prime},\mathbb{Q}^{\prime}$ such that $\operatorname{TV}(\mathbb{P},\mathbb{P}^{\prime}),\operatorname{TV}(\mathbb{Q},\mathbb{Q}^{\prime})=o(1)$, then there exists a constant $C$ such that no algorithm with running time $n^{D/(\log n)^{C}}$ that achieves strong detection between $\mathbb{P}$ and $\mathbb{Q}$.
@@
    If $\mathsf{Adv}_{\leq D}\big{(}\tfrac{\mathrm{d}\mathbb{P}^{\prime}}{\mathrm{d}\mathbb{Q}^{\prime}}\big{)}=1+o(1)$ as $n\to\infty$ for some $\mathbb{P}^{\prime},\mathbb{Q}^{\prime}$ such that $\operatorname{TV}(\mathbb{P},\mathbb{P}^{\prime}),\operatorname{TV}(\mathbb{Q},\mathbb{Q}^{\prime})=o(1)$, then there exists a constant $C$ such that no algorithm with running time $n^{D/(\log n)^{C}}$ that achieves weak detection between $\mathbb{P}$ and $\mathbb{Q}$.
@@
Assuming Conjecture C.1, there exists a constant $C$ such that for all $1\leq i\leq n$ and all statistics $\{g_{i,j}=g_{i,j}(A,B):1\leq j\leq n\}$ such that each $g_{i,j}(A,B)$ can be computed in time $n^{D/(\log n)^{C}}$ and
@@
Without losing of generality, we may assume that $i=1$ in the following proof. Suppose on the contrary that there are statistics $\{g_{j}=g_{j}(A,B):1\leq j\leq n\}$ that can be computed in time $n^{D/(\log n)^{C}}$ such that $\sum_{1\leq j\leq n}g_{j}(A,B)=o(n)$ for all graphs $A,B$ and (3.6) holds. Again we may assume that $0\leq g_{i}\leq 1$. Recall (3.7), (3.8) and (3.9). Using Lemma 3.5 and Theorem 2.4, we see that (assuming Conjecture C.1) we have (3.10) still holds. To this end, define $g=g_{1}+\ldots+g_{n}$. Using $0\leq g\leq n$, we see that (3.10) yields that $\mathbb{Q}(g\geq\Omega(n))\geq\Omega(n)$, which violates our assumption (C.1). This implies that (3.6) is impossible and thus completes our proof. ∎
@@
Suppose on the contrary there are statistics $\{f_{j}:1\leq j\leq n\}$ that can be computed in time $n^{D/(\log n)^{C}}$ satisfying Items (1) and (2) in Definition 3.1 with $\mathbb{E}_{\mathbb{P}_{*}}[f_{\pi_{*}(i)}]\geq 1-c$ for some fixed constant $0.010<c<$ 0.01. We first claim that we may assume that